# Predictive Coherence and the Moment Hierarchy: Martingale Posteriors for Exchangeable Bernoulli Sequences

### 05 Mar 2026 v2

@@ 4.1 Moment representation
Theorem 4.1 (Moment representation).
@@
By Theorem 3.2, $X_{n+1},\ldots,X_{n+k}$ are conditionally i.i.d. $\mathrm{Bern}(\theta)$ and conditionally independent of $\mathcal{F}_{n}$. Hence $\mathbb{P}(X_{n+1:n+k}=x_{1:k}\mid\mathcal{F}_{n},\theta)=\theta^{s}(1-\theta)^{k-s}$. Integrate over $\Pi(\,\mathrm{d}\theta\mid\mathcal{F}_{n})$ by the tower property. ∎
Corollary 4.2 (Moment expansion).
@@ 4.2 Injectivity
By Corollary 4.2, knowing $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ for all $k\geq 1$ determines the moment sequence $(\mathbb{E}[\theta^{j}\mid\mathcal{F}_{n}])_{j\geq 1}$ via Möbius inversion (Proposition 4.6). On $[0,1]$, the moment sequence uniquely determines the measure (Hausdorff moment theorem; Theorem A.1 in the Appendix). ∎
Remark 4.4 (Polynomial density).
Theorem 4.3 also follows from Weierstrass approximation: polynomials are dense in $C[0,1]$, so the moment sequence determines all expectations $\mathbb{E}[f(\theta)\mid\mathcal{F}_{n}]$ for $f\in C[0,1]$, hence the measure. Injectivity relies on the compact support $[0,1]$; on $\mathbb{R}$, the Stieltjes moment problem may have non-unique solutions (e.g., the log-normal distribution).
Remark 4.5 (Comparison with the real line).
On $[0,1]$, the moment problem is determinate: a probability measure is uniquely determined by its moment sequence (Theorem A.1). This is the mechanism behind Theorem 4.3. On $\mathbb{R}$, the Hamburger moment problem is indeterminate in general: distinct measures can share all moments. A classical example is the log-normal: $X\sim\mathrm{LogNormal}(0,1)$ has the same moments as the family $X\cdot(1+a\sin(2\pi\log X))$ for $|a|\leq 1$. The moment hierarchy established in Section 4 therefore provides a complete characterization of the posterior only because $\theta$ takes values in the compact interval $[0,1]$. For parameters in $\mathbb{R}^{d}$, analogous results would require additional conditions such as exponential moment bounds (Carleman’s condition) or restriction to exponential families with compact sufficient statistics.
@@ 4.3 Binomial inversion
Proposition 4.6 (Möbius inversion).
Let $m_{k}=\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ and $\mu_{j}=\mathbb{E}[\theta^{j}\mid\mathcal{F}_{n}]$ with $\mu_{0}=1$. Then
$\mu_{j}=\sum_{\ell=0}^{j}\binom{j}{\ell}(-1)^{\ell}\,m_{\ell},\qquad j\geq 0.$ (11)
The matrix relating $(\mu_{0},\ldots,\mu_{k})$ to $(m_{0},\ldots,m_{k})$ is upper-triangular with diagonal entries $\pm 1$, hence invertible.
@@
Example 4.7 (Inversion at $j=2$).
From the forward relation: $m_{2}=1-2\mu_{1}+\mu_{2}$. By Möbius inversion (11): $\mu_{2}=m_{0}-2m_{1}+m_{2}$, i.e., $\mu_{2}=1-2(1-\mu_{1})+m_{2}=2\mu_{1}-1+m_{2}$. Substituting $m_{2}=(1-\mu_{1})^{2}+\sigma_{n}^{2}$ gives $\mu_{2}=2\mu_{1}-1+(1-\mu_{1})^{2}+\sigma_{n}^{2}=\mu_{1}^{2}+\sigma_{n}^{2}$, recovering $\sigma_{n}^{2}=\mu_{2}-\mu_{1}^{2}$ as expected. This illustrates the triangular structure: the $j$-th moment and the $j$-step run probability determine each other given lower-order quantities.
@@ 4.4 Explicit expansion for small $k$
5 KL geometry and the moment hierarchy
This section develops the KL geometry underlying the moment hierarchy, adding two structural results (Theorem 5.1 and Proposition 5.2) and a diagram.
For Bernoulli, $D_{\mathrm{KL}}(p\|\cdot)$ has Fisher information $\mathcal{I}(\theta)=1/(\theta(1-\theta))$. Since $m_{n}$ minimises $D_{\mathrm{KL}}(m_{n}\|\cdot)$, its gradient vanishes at $m_{n}$; the curvature there is $\mathcal{I}(m_{n})=1/(m_{n}(1-m_{n}))$. The posterior variance satisfies $\sigma_{n}^{2}\approx m_{n}(1-m_{n})/n$ (Bernstein–von Mises), the reciprocal of this curvature. Predictively, the hierarchy mirrors a derivative hierarchy:
@@ 5 KL geometry and the moment hierarchy
5.1 The Sanov Gap
Theorem 5.1 (Sanov Gap).
@@ 5.1 The Sanov Gap
    Non-pathological gap. For any $m\in(0,1)$ and $k\geq 2$, the $k$-step predictive ranges over a nondegenerate interval as $\nu$ ranges over posteriors with mean $m$; the gap is not a boundary phenomenon (Remark 6.6).
@@
(i) $D_{\mathrm{KL}}(m\|\theta)$ is minimised at $\theta=m$, so its derivative vanishes there (Theorem 2.1(ii)). (ii) is equation (15) (Proposition 6.4). (iii) follows from Lemma 6.1 and Remark $6.$ ∎
Theorem 5.1 is the geometric source of the non-uniqueness in Theorem 6.3: posteriors sharing the same mean are first-order equivalent in the Sanov expansion but separate at second order via $\sigma^{2}$.
5.2 Moderate deviations and the second moment as first nontrivial Sanov term
Proposition 5.2 (Second moment as leading Sanov curvature).
Under the moderate-deviations scaling ($a_{n}/\sqrt{n}\to\infty$, $a_{n}=o(\sqrt{n})$), the posterior satisfies a moderate-deviation principle with quadratic rate $\frac{1}{2}(\theta-\theta_{0})^{2}\mathcal{I}(\theta_{0})$ (Eichelsbacher and Ganesh, 2002). Consequently:
  1. (i)
    The posterior variance satisfies $\sigma_{n}^{2}=m_{n}(1-m_{n})/n+O(n^{-2})$, recovering the Bernstein–von Mises rate.
  2. (ii)
    The second moment is the first nontrivial curvature term: it corresponds to $\nabla^{2}D_{\mathrm{KL}}|_{m_{n}}$ and cannot be recovered from the zeroth- and first-order terms.
  3. (iii)
    Specifying only the posterior mean (first-moment coherence) is equivalent to retaining the zeroth-order Sanov term; the second moment requires the quadratic term; the $k$-th moment requires the $k$-th derivative.
@@ 6.1 Non-identification and moment insufficiency
Remark 6.2 (Extension to general functionals).
Lemma 6.1 extends to any non-affine continuous $f:[0,1]\to\mathbb{R}$: the mean does not determine $\mathbb{E}[f(\theta)\mid\mathcal{F}_{n}]$ (Appendix B).
@@ 6.2 Numerical illustration
Theorem 6.3 (Moment insufficiency).
@@
By Lemma 6.1 and Theorem 4.1: any two measures $\nu_{1},\nu_{2}$ on $[0,1]$ with equal mean but different $k$-th moments (for $k\geq 2$) serve as de Finetti mixing measures with the same posterior mean $m_{n}$ but different $k$-step predictives. ∎
@@ 6.3 Quantitative bound
Proposition 6.4 (Second-order discrepancy).
@@
Set $f(\theta)=(1-\theta)^{k}$. By Taylor’s theorem with Lagrange remainder around $m_{n}$: $f(\theta)=f(m_{n})+f^{\prime}(m_{n})(\theta-m_{n})+\tfrac{1}{2}f^{\prime\prime}(\xi)(\theta-m_{n})^{2}$. Take conditional expectations and use $\mathbb{E}[\theta-m_{n}\mid\mathcal{F}_{n}]=0$:
$\mathbb{E}[f(\theta)\mid\mathcal{F}_{n}]-f(m_{n})=\tfrac{1}{2}\mathbb{E}[f^{\prime\prime}(\xi)(\theta-m_{n})^{2}\mid\mathcal{F}_{n}].$
Since $f^{\prime\prime}(\theta)=k(k-1)(1-\theta)^{k-2}\geq 0$ on $[0,1]$ (strict convexity) and $\sup f^{\prime\prime}=k(k-1)$, the mean value theorem for integrals gives (15). The lower bound is strict when $\sigma_{n}^{2}>0$ by strict convexity. ∎
Corollary 6.5 (Plug-in as Dirac).
@@
Remark 6.6 (Strength of non-identification).
The non-identification in Lemma 6.1 is not a pathological boundary phenomenon. For any $m\in(0,1)$ and any $k\geq 2$, the set of posteriors with mean $m$ is an infinite-dimensional convex set, and the $k$-step predictive $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ takes all values in the interval $\bigl[(1-m)^{k},\,(1-m)^{k}+\tfrac{k(k-1)}{2}m(1-m)\bigr)$ as the posterior ranges over distributions with mean $m$. The lower bound is attained by $\delta_{m}$ (plug-in) and the upper bound is approached by two-point distributions concentrating near 0 and $1$. At $k=2$ with $m=1/2$, the Bayes predictive ranges from $1/4$ (degenerate) to $1/2$ (symmetric Bernoulli on $\{0,1\}$), a factor of two.
Remark 6.7 (Scope).
Theorem 6.3 does not assert that the martingale posterior framework is incompatible with Bayesian inference. Every Bayesian posterior mean is a martingale (Remark 3.6). The theorem asserts that condition (1) alone does not determine the full posterior. A specific likelihood, a named prior, or explicit higher-moment constraints are needed for multi-step predictive coherence (Definition 3.7).
@@ 7 Decision-theoretic consequence
Let $\delta$ be a predictive rule producing $\hat{P}_{\delta}(\cdot\mid\mathcal{F}_{n})$ on $\{0,1\}^{k}$.
Definition 7.1.

### 01 Apr 2026 v5

@@ 4.1 Moment representation
Theorem 13 (Moment representation).
@@
By Theorem 3, $X_{n+1},\ldots,X_{n+k}$ are conditionally i.i.d. $\mathrm{Bern}(\theta)$ and conditionally independent of $\mathcal{F}_{n}$. Hence $\mathbb{P}(X_{n+1:n+k}=x_{1:k}\mid\mathcal{F}_{n},\theta)=\theta^{s}(1-\theta)^{k-s}$. Integrate over $\Pi(\,\mathrm{d}\theta\mid\mathcal{F}_{n})$ by the tower property. ∎
Corollary 14 (Moment expansion).
@@ 4.2 Injectivity
The argument proceeds in two steps.
Step 1: run probabilities determine moments. By Möbius inversion (Proposition 18), $\mathbb{E}[\theta^{j}\mid\mathcal{F}_{n}]=\sum_{\ell=0}^{j}\binom{j}{\ell}(-1)^{\ell}\,\mathbb{E}[(1-\theta)^{\ell}\mid\mathcal{F}_{n}]$ for each $j\geq 1$. The matrix relating the run-probability sequence $(r_{0},r_{1},\ldots)$ to the moment sequence $(\mu_{0},\mu_{1},\ldots)$ is upper-triangular with $\pm 1$ on the diagonal, hence invertible. Knowing $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ for all $k\geq 1$ therefore determines $\mathbb{E}[\theta^{j}\mid\mathcal{F}_{n}]$ for all $j\geq 1$.
Step 2: moments determine the measure on $[0,1]$. Since $\theta\in[0,1]$, the Hausdorff moment theorem (Theorem 51) guarantees that the moment sequence $(\mathbb{E}[\theta^{j}\mid\mathcal{F}_{n}])_{j\geq 0}$ uniquely determines the conditional law $\Pi(\cdot\mid\mathcal{F}_{n})$. Combining Steps 1 and 2 establishes injectivity. ∎
Remark 16 (Polynomial density).
Theorem 15 also follows from Weierstrass approximation: polynomials are dense in $C[0,1]$, so the moment sequence determines all expectations $\mathbb{E}[f(\theta)\mid\mathcal{F}_{n}]$ for $f\in C[0,1]$, hence the measure. Injectivity relies on the compact support $[0,1]$; on $\mathbb{R}$, the Stieltjes moment problem may have non-unique solutions (e.g., the log-normal distribution).
Remark 17 (Comparison with the real line).
On $[0,1]$, the moment problem is determinate: a probability measure is uniquely determined by its moment sequence (Theorem 51). This is the mechanism behind Theorem 15. On $\mathbb{R}$, the Hamburger moment problem is indeterminate in general: distinct measures can share all moments. A classical example is the log-normal: $X\sim\mathrm{LogNormal}(0,1)$ has the same moments as the family $X\cdot(1+a\sin(2\pi\log X))$ for $|a|\leq 1$. The moment hierarchy established in Section 4 therefore provides a complete characterization of the posterior only because $\theta$ takes values in the compact interval $[0,1]$. For parameters in $\mathbb{R}^{d}$, analogous results would require additional conditions such as exponential moment bounds (Carleman’s condition) or restriction to exponential families with compact sufficient statistics.
@@ 4.3 Binomial inversion
Proposition 18 (Möbius inversion).
Let $r_{k}=\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ denote the $k$-step run probability and $\mu_{j}=\mathbb{E}[\theta^{j}\mid\mathcal{F}_{n}]$ with $\mu_{0}=1$. Then
$\mu_{j}=\sum_{\ell=0}^{j}\binom{j}{\ell}(-1)^{\ell}\,r_{\ell},\qquad j\geq 0.$ (10)
The matrix relating $(\mu_{0},\ldots,\mu_{k})$ to $(r_{0},\ldots,r_{k})$ is upper-triangular with diagonal entries $\pm 1$, hence invertible.
@@
Example 19 (Inversion at $j=2$).
From the forward relation: $r_{2}=1-2\mu_{1}+\mu_{2}$. By Möbius inversion (10): $\mu_{2}=r_{0}-2r_{1}+r_{2}$, i.e., $\mu_{2}=1-2(1-\mu_{1})+r_{2}=2\mu_{1}-1+r_{2}$. Substituting $r_{2}=(1-\mu_{1})^{2}+\sigma_{n}^{2}$ gives $\mu_{2}=2\mu_{1}-1+(1-\mu_{1})^{2}+\sigma_{n}^{2}=\mu_{1}^{2}+\sigma_{n}^{2}$, recovering $\sigma_{n}^{2}=\mu_{2}-\mu_{1}^{2}$ as expected. This illustrates the triangular structure: the $j$-th moment and the $j$-step run probability determine each other given lower-order quantities.
@@ 4.4 Explicit expansion for small $k$
5 KL geometric interpretation
This section gives a geometric reading of the moment hierarchy in terms of the Sanov rate function. It adds one result (Theorem 20) and a figure; it is not required for subsequent sections.
For Bernoulli, $D_{\mathrm{KL}}(p\|\cdot)$ has Fisher information $\mathcal{I}(\theta)=1/(\theta(1-\theta))$. For any fixed $q\in(0,1)$, the function $\theta\mapsto D_{\mathrm{KL}}(q\|\theta)$ is minimised at $\theta=q$, where its first derivative vanishes and its second derivative equals $\mathcal{I}(q)=1/(q(1-q))$. Evaluating at $q=m_{n}$, the curvature is $\mathcal{I}(m_{n})=1/(m_{n}(1-m_{n}))$. This is a property of the KL divergence, not of the posterior: it should not be confused with the posterior mean’s role in Bayesian updating. The posterior variance satisfies $\sigma_{n}^{2}\approx m_{n}(1-m_{n})/n$ (Bernstein–von Mises), the reciprocal of this curvature. Predictively, the hierarchy mirrors a derivative hierarchy:
@@ 5 KL geometric interpretation
5.1 First-order indistinguishability and second-order separation
Theorem 20 (Geometric separation).
@@ 5.1 First-order indistinguishability and second-order separation
    Non-pathological gap. For any $m\in(0,1)$ and $k\geq 2$, the $k$-step predictive ranges over a nondegenerate interval as $\nu$ ranges over posteriors with mean $m$; the gap is not a boundary phenomenon (Remark 26).
@@
(i) The Sanov rate function $D_{\mathrm{KL}}(m\|\theta)$ achieves its minimum at $\theta=m$ (Theorem 1(ii)), so the first derivative vanishes: $\partial_{\theta}D_{\mathrm{KL}}(m\|\theta)\big|_{\theta=m}=0$. Any two posteriors $\nu_{1},\nu_{2}$ sharing mean $m$ integrate this linear term to the same value, producing identical first-order contributions to the posterior log-likelihood.
(ii) The second derivative of $f(\theta)=(1-\theta)^{k}$ at $\theta=m$ is $k(k-1)(1-m)^{k-2}>0$ for $k\geq 2$ and $m\in(0,1)$. By Proposition 24, the $k$-step predictive satisfies $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m)^{k}=\frac{k(k-1)}{2}(1-\xi)^{k-2}\sigma^{2}$, with the correction proportional to $\sigma^{2}$. Distinct posteriors with distinct variances therefore yield distinct $k$-step predictives.
(iii) By Lemma 21, for any $m\in(0,1)$ and $k\geq 2$, the point mass $\delta_{m}$ and a two-point distribution with mean $m$ achieve different values of $\mathbb{E}[(1-\theta)^{k}]$. Remark 26 shows the achievable range is a nondegenerate interval, so the gap is not a boundary artefact. ∎
Geometrically, posteriors sharing the same mean are first-order equivalent in the Sanov expansion but separate at second order via $\sigma^{2}$; this is the mechanism underlying Theorem 23.
5.2 Geometric interpretation of the moment hierarchy
Under standard Bernstein–von Mises conditions, the posterior variance is of order $1/n$, and the quadratic term in the KL expansion corresponds heuristically to the leading curvature contribution. The second moment is the first nontrivial curvature term: it corresponds to $\nabla^{2}D_{\mathrm{KL}}|_{m_{n}}$ and cannot be recovered from the zeroth- and first-order terms. Specifying only the posterior mean retains the approximate location of the KL bowl; the second moment requires the quadratic curvature; the $k$-th moment requires the $k$-th derivative.
@@ 6.1 Non-identification and moment insufficiency
Remark 22 (Extension to general functionals).
Lemma 21 extends to any non-affine continuous $f:[0,1]\to\mathbb{R}$: the mean does not determine $\mathbb{E}[f(\theta)\mid\mathcal{F}_{n}]$ (Appendix B).
@@ 6.2 Numerical illustration
Theorem 23 (Moment insufficiency).
@@
Fix $m_{n}\in(0,1)$ and $k\geq 2$. We construct two posteriors sharing mean $m_{n}$ but yielding distinct $k$-step predictives.
Let $\nu_{1}=\delta_{m_{n}}$, so $\int(1-\theta)^{k}\,\mathrm{d}\nu_{1}=(1-m_{n})^{k}$.
Choose $0<a<m_{n}<b<1$ and set $\nu_{2}=\lambda\delta_{a}+(1-\lambda)\delta_{b}$ with $\lambda=(b-m_{n})/(b-a)$. Then $\int\theta\,\mathrm{d}\nu_{2}=m_{n}$ by construction, but $\nu_{2}$ has variance $\sigma^{2}=\lambda(1-\lambda)(b-a)^{2}>0$. Since $\theta\mapsto(1-\theta)^{k}$ is strictly convex for $k\geq 2$, Jensen’s inequality gives $\int(1-\theta)^{k}\,\mathrm{d}\nu_{2}>(1-m_{n})^{k}=\int(1-\theta)^{k}\,\mathrm{d}\nu_{1}$.
By Theorem 13, the $k$-step run probability under posterior $\nu_{i}$ is $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]=\int(1-\theta)^{k}\,\mathrm{d}\nu_{i}$, so the two posteriors yield distinct $k$-step predictives despite sharing the same mean. Since the choice of $a,b$ is arbitrary (subject to $0<a<m_{n}<b<1$), the mapping $m_{n}\mapsto\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ is set-valued. ∎
@@ 6.3 Quantitative bound
Proposition 24 (Second-order discrepancy).
@@
Set $f(\theta)=(1-\theta)^{k}$. By Taylor’s theorem with Lagrange remainder, for each $\theta\in[0,1]$ there exists $\xi(\theta)\in[\min(\theta,m_{n}),\,\max(\theta,m_{n})]$ such that $f(\theta)=f(m_{n})+f^{\prime}(m_{n})(\theta-m_{n})+\tfrac{1}{2}f^{\prime\prime}(\xi(\theta))(\theta-m_{n})^{2}$. Take conditional expectations. Since $\mathbb{E}[\theta-m_{n}\mid\mathcal{F}_{n}]=0$, the linear term vanishes:
$\mathbb{E}[f(\theta)\mid\mathcal{F}_{n}]-f(m_{n})=\tfrac{1}{2}\,\mathbb{E}\bigl[f^{\prime\prime}(\xi(\theta))\,(\theta-m_{n})^{2}\mid\mathcal{F}_{n}\bigr].$
Now $f^{\prime\prime}(\theta)=k(k-1)(1-\theta)^{k-2}$ is continuous and nonnegative on $[0,1]$, with $0\leq f^{\prime\prime}(\theta)\leq k(k-1)$. Since $f^{\prime\prime}(\xi(\theta))(\theta-m_{n})^{2}\geq 0$, the integral on the right is nonnegative. By the mean value theorem for integrals (applied to the probability measure $(\theta-m_{n})^{2}\,\mathrm{d}\Pi(\theta\mid\mathcal{F}_{n})/\sigma_{n}^{2}$ when $\sigma_{n}^{2}>0$), there exists $\xi_{n}\in[0,1]$ such that
$\mathbb{E}\bigl[f^{\prime\prime}(\xi(\theta))\,(\theta-m_{n})^{2}\mid\mathcal{F}_{n}\bigr]=f^{\prime\prime}(\xi_{n})\,\sigma_{n}^{2}=k(k-1)(1-\xi_{n})^{k-2}\,\sigma_{n}^{2}.$
This gives (14). The upper bound (15) follows from $(1-\xi_{n})^{k-2}\leq 1$. The lower bound in (15) is strict when $\sigma_{n}^{2}>0$ because $f^{\prime\prime}>0$ on $(0,1)$. ∎
Corollary 25 (Plug-in as Dirac).
@@
Remark 26 (Strength of non-identification).
The non-identification in Lemma 21 is not a pathological boundary phenomenon. For any $m\in(0,1)$ and any $k\geq 2$, as the posterior ranges over all distributions with mean $m$, the $k$-step predictive $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ takes all values in the interval $\bigl[(1-m)^{k},\;1-m\bigr]$. The lower bound $(1-m)^{k}$ is attained by $\delta_{m}$ (the plug-in value, via Jensen’s inequality and strict convexity of $(1-\theta)^{k}$). The upper bound 1m1-m is attained by the two-point distribution $(1-m)\delta_{0}+m\delta_{1}$, which has mean $m$ and gives $\mathbb{E}[(1-\theta)^{k}]=(1-m)\cdot 1^{k}+m\cdot 0^{k}=1-m$. At $k=2$ with $m=1/2$, the predictive ranges from $1/4$ (degenerate) to $1/2$ (endpoint distribution), a factor of two.
Figure 2: Moment insufficiency: for fixed posterior mean $m$, the $k$-step predictive $\mathbb{E}[(1-\theta)^{k}]$ ranges over the shaded region $\bigl[(1-m)^{k},\;1-m\bigr]$ (Remark 26). Dashed: degenerate lower bound $(1-m)^{k}$ (plug-in). Dotted: two-point upper bound 1m1-m. Curves show Beta(cm,c(1m))(cm,\,c(1-m)) posteriors at concentrations $c\in\{1,2,5,20,100\}$. Higher concentration drives the predictive toward the degenerate plug-in; lower concentration spreads posterior mass and increases the run probability. All values are computed exactly via $\mathbb{E}[(1-\theta)^{k}]=(b)_{k}/(a+b)_{k}$.
Remark 27 (Scope).
Theorem 23 does not assert that the martingale posterior framework is incompatible with Bayesian inference. Every Bayesian posterior mean is a martingale (Remark 7). The theorem asserts that condition (1) alone does not determine the full posterior. A specific likelihood, a named prior, or explicit higher-moment constraints are needed for multi-step predictive coherence (Definition 8).
@@ 7 Decision-theoretic consequence
The insufficiency theorem has an immediate decision-theoretic corollary: under any strictly proper scoring rule, the plug-in predictive is inadmissible. Let $\delta$ be a predictive rule producing $\hat{P}_{\delta}(\cdot\mid\mathcal{F}_{n})$ on $\{0,1\}^{k}$.
Definition 28.